# A quantitative analysis of semantic information in deep representations of text and images

### v3

@@ 3.1.1 Heterogeneity of DeepSeek’s representations for different language pairs
We repeated the previous translation experiment with other language pairs, in particular English-Italian, English-Spanish, Italian-Spanish and German-Hungarian. We considered again sentences taken from opus_books opus_books and applied the same filtering procedure regarding their number of tokens, as in Fig. 1a).
Fig. 1b) shows the same qualitative behavior observed for English-Spanish pairs (low information imbalance in the deep layers) but with major quantitative differences. Even if the training dataset of DeepSeek-V3 contains several languages in which the network can perform well, the quality of the representations is heterogeneous across languages, since it depends on the amount of data seen in each language. Fig. 1b) shows that the information imbalance between English and Italian is higher than between English and Spanish. We conjecture that this difference is due to the greater online presence of the Spanish language relative to Italian LanguagesUsedInternet2025 , and the consequent difference in the amount of training data. Aligned with this argument, we observe that, even though Spanish and Italian are linguistically closer than either is to English, the Information Imbalance between Spanish and Italian is higher than between English-Spanish or English-Italian translation pairs. Finally, we observe that translations involving a less common language, namely Hungarian, are associated with a significantly higher II (the minimum value is 0.5).
Figure 2: Panel a) Minimum Information Imbalance (II) across depth, from English to Spanish representations, as a function of the number of tokens used in the computation, for DeepSeek-V3 and Llama3.1-8b, when considering shorter (40 to 80 tokens, in circles) or longer (100 to 200 tokens, in triangles) sentences. Panels b) and c) II from the last token to a previous token at token-distance $\tau$, using English sentences, computed on the representations generated by DeepSeek-V3 in panel a) and by Llama3.1-8b in panel b), in different layers. Long-distance token-token correlations are maximal (II increases most slowly) for representations on the inner semantic layers (43 for DeepSeek-V3, 23 for Llama3.1-8b), and the effect is dramatically stronger in DeepSeek-V3. Error bars computed with 5 repetitions of a Jackknife procedure, subsampling $2000$ sentences out of $5000$, are smaller than marker size.
@@ 3.1.2 Semantic information in LLMs is spread across many tokens
Having shown that long-span Information Imbalance captures the semantic similarity of sentences in different languages, we turn our attention to the extent to which different tokens are contributing to cross-linguistic similarity. Fig. 2a) shows the minimum II across all layers as a function of the number of tokens involved in the calculation, again for English-Spanish translations from opus_books. The span is determined starting from the end and moving back towards the sentence beginning. The main result here, for both models, is that the addition of more tokens significantly increases the mutual predictivity of the translation representations, proving that semantic information is not concentrated in the last tokens, but spread over many of them.
Focusing first on DeepSeek-V3, we note that, up to about 10 tokens, the results do not change much between shorter (40 to 80 tokens) and longer (100 to 200 tokens) sentences. In both cases adding more tokens significantly boosts the information content. However, for shorter sentences we observe that the II reaches its lowest value at around $20$ tokens, and it increases with the addition of further tokens. We conjecture that this is because the last added tokens are close to the start of the sentence, and thus they do not carry much information about the semantic content of the whole sentence, given the autoregressive nature of the attention mechanism in the considered language models. In contrast, for longer sentences, we observe that adding tokens leads to a rapid decrease in II, which converges to very low values (around $0.05$) at around 50 tokens and stays there. On the one hand, this reflects the fact that, as longer sentences contain more information, their translation will contain more shared information at the semantic level. On the other hand, it further confirms that information is spread across a large number of tokens. We observe similar trends for Llama3.1-8b II, although the curves are systematically above the DeepSeek-V3 ones. Similar to what we observed in Fig 1a), this confirms that bigger and better performing models generate representations of better quality, capturing deep semantic similarity across different languages. Finally, this result shows that not every token contributes homogeneously to the semantic similarity between representations, and thus this analysis suggests to take the concatenation of roughly the last 20 tokens as representative of the sentence content.
@@ 3.1.3 Inner representations have longer-range correlations
The previous analysis implies that the representation of a sentence in the deep layers, where it is transferrable between different languages, is encoded in several tokens. One wonders if the activations of the tokens in those layers are uncorrelated or not. If activations were independent, one could argue that different tokens are associated to different “directions” in semantic space. If they are correlated, this would suggest that the model is building a joint semantic representation of sentences spreading its information across its component tokens. We address this question by measuring how informative two distinct tokens of the same sentence are about each other. Fig. 2 (panels b) and c)) shows how much the II between two token positions depends on their distance, for sentences in English. Concretely, we measure $\Delta(T\rightarrow T-\tau)$, the II between the token at position $T$, and tokens at position $T-\tau$, for $\tau=1,2,$.... We report this measure for the first and last layers, as well as an intermediate layer from the semantic region (layers 43 of DeepSeek-V3 and 23 of Llama3.1-8b, respectively). For both LLMs, the II rapidly grows as a function of $\tau$ for the initial layers and saturates to 1, clearly showing that the first layer is short-range correlated. For the deeper layers that were shown in Fig. 1 to correspond to the phase of shared cross-linguistic information, the II exhibits the slowest growth, meaning that far away tokens are predictive about each other, i.e., in these layers the information is mostly shared between tokens. DeepSeek’s deep representations have notably lower II across tokens than Llama3.1-8b’s, making it evident that better-quality semantic representations more clearly capture long-distance correlations. A clear separation in effect strength between the intermediate semantic layer and the last layer (61 for DeepSeek-V3, 32 for Llama3.1-8b) is moreover only present for DeepSeek-V3. Still, even for DeepSeek-V3 the last layer has II values clearly below those of the first layer, suggesting that, while the last layer is optimized to write the output, it still carries semantic content.
Figure 3: Panel a) In solid markers, the “backward” Information Imbalance (II), i.e., from the last token to a previous token at token-distance $\tau$, as a function of $\tau$, for DeepSeek-V3 representations at layer 43 and sentences in several languages. In dashed lines, the reciprocal “forward” II. Panel b) Asymmetry $\mathcal{A}$ between forward and backward II, averaged over $\tau$, as a function of DeepSeek’s relative depth. Error bars computed with 5 repetitions of a Jackknife procedure, subsampling $1000$ sentences out of $2000$, are smaller than marker size.
3.1.4 Correlations between tokens as a hallmark of quality in semantic representations
Fig. 3a) focuses on a semantic layer from DeepSeek-V3, and it shows token-token II as a function of token distance for several languages. We present the II in both directions: The "backward" II, from the last token, $T$, to a previous token at distance $\tau$, $\Delta(T\rightarrow T-\tau)$, and the reciprocal or “forward” II, $\Delta(T-\tau\rightarrow T)$. We observe that English deep representations are much more correlated than the representations in the other languages. We conjecture that this is a sign of better quality, consistent with Fig. 1b), where the presence of English in a pair results in the best predictability (lowest II scores). At the opposite extreme, note that single-language German and Hungarian representations are the least correlated in Fig. 3a), and the German-Hungarian translations have indeed the worst mutual predictability score (highest II) in Fig. 1. Furthermore, Spanish and Italian have the second and third most correlated inner representations in Fig. 3a), coherently with the profiles of the English-Spanish and English-Italian pairs in Fig. 1b) (lowest and second lowest IIs, respectively).
Finally, Fig. 3a) shows that, for English, as the distance between tokens $\tau$ increases, the relative information content between past and future tokens gets increasingly asymmetric. In particular, earlier tokens are more informative of the last one than the other way around. This could be expected, since causal models like DeepSeek-V3 are trained to predict future tokens, and text itself has a natural causal asymmetry. Nonetheless, Fig. 3a) shows that this past-to-future asymmetry is heterogeneous across languages. To further study the effect, we first define the asymmetry as $\mathcal{A}=\Delta(T\rightarrow T-\tau)-\Delta(T-\tau\rightarrow T)$. When $\mathcal{A}$ is positive, the earlier token predicts the last one more than the latter predicts the earlier token, whereas a negative $\mathcal{A}$ has the opposite interpretation. Fig. 3b) shows the information asymmetry $\mathcal{A}$ of DeepSeek-V3 as a function of its layers for the all languages we studied. For initial layers (i.e., relative depth less than roughly $0.4$, see Fig. 1), we observe that all representations have a comparable amount of causal asymmetry. Instead, for semantic layers there are strong heterogeneities between languages. English remains causally asymmetric until the last few layers of the network, while the other languages present earlier negative values of $\mathcal{A}$. Intriguingly, Spanish and Italian, the two languages with the lowest IIs from English, have a later dip in $\mathcal{A}$ compared to German and Hungarian. These results call for further experiments to study possible relationships between these observed information asymmetries, model performance, and linguistic structure. We leave them as material for future research.
@@ 3.2 Images sharing high level semantics
We extend the analysis to pairs of images and measure how vision transformers (ViTs) capture semantic information for two systems: image-gpt pmlr-v119-chen20s and DinoV2 oquabDINOv2LearningRobust2024 . Image-gpt performs next-token prediction after it (i) down-scales the images; (ii) unrolls them, juxtaposing the rows; (iii) quantizes the colors. It mimics the strategy of LLMs: the last layers must perform next-token prediction; previous work suggests that semantics appears in the middle pmlr-v119-chen20s . DinoV2 encodes salient features. It trains a student network to mimic its teacher and reconstruct different instances of the same image. The final layer should be the most semantically rich: its output is the input to downstream tasks such as depth estimation, image segmentation, and instance retrieval.
We first compare representations for pairs of images from the Imagenet1k dataset Deng:etal:2009 . Each couple shares the same class, which we take as a proxy of similar semantics: we process the pairs of instances with the same architecture and compute the information imbalance on a random subset of 2500 same-class pairs of images, which we sample without repetition. We replicate the procedure five times to quantify uncertainty. We report the results in the left panel of Fig. 4. As the assignment of same-pair images to two spaces used to compute II is arbitrary, we report the average of II in the two directions. The minimum of the information imbalance is in the middle for image-gpt and at the end for DinoV2, consistently with previous work valeriani2023geometry . The information imbalance is lower for DinoV2 ($\approx 0.43$): this would correspond to roughly 10% shared features in the Gaussian case, see Supp. Inf., Sec. C.
@@ 3.3 Images and captions: multimodal data sharing semantic content
Image-caption pairs encode the same semantic content in different modalities: we are interested in their representations as semantics must play a key role in alignment. We process text with DeepSeek-V3 and images with DinoV2 and image-gpt. The dataset is flickr30k youngImageDescriptionsVisual2014 : it contains pairs of images and captions; the captions are the concatenations of five different human-generated descriptions. We next repeat the analysis of Fig. 2a), estimating the minimum Information Imbalance between the image transformer and DeepSeek-V3 as a function of the number of image tokens concatenated from the end of the sequence. For DeepSeekV3, we fix the token window to 9, which corresponds to the point where the steep decrease of II observed in Fig. 2a) essentially saturates. We find that, for both DinoV2 and image-gpt, it is necessary to concatenate the last 200 tokens to capture the relevant shared information. Results obtained with shorter windows (last 50 tokens) and with mean-token are reported for comparison in Supp. Inf., Sec. J.
@@
In Fig. 4 we report the information imbalance between layer 52 of DeepSeek-V3 (captions), DinoV2 and image-gpt (images). We select layer 52 of DeepSeek-V3 because it is a layer in which the representations of text in different languages is highly mutually predictive, see Fig. 1. We recover the exact same semantic regions provided by the analysis of the image-image pairs: towards the end for DinoV2 and in the middle for image-gpt. DinoV2 has a lower information imbalance of 0.4. Using the Gaussian process as a reference (Supp. Inf., Sec. C), this would correspond to approximately 10% of shared features. The minimum II for image-gpt is $\approx$ 0.6, corresponding to 5% of shared features. The dashed lines in Fig. 4 report the information imbalance from the image to the text representation. Remarkably, for the image-gpt model the II is considerably larger than in the text-to-image direction. In particular, the minimum value of the II is $\approx 0.7$. The difference between the II in the two direction is significantly higher than the error bars estimated by bootstrap, indicating that the effect is statistically significant. This difference is likely due to the better quality of the DeepSeek-V3 representations. The asymmetry in the information imbalance is reversed for the DinoV2 model. In this case the image representation predicts the text representation marginally better than the reverse. This can be due to the fact that DinoV2 is explicitly trained to extract semantically meaningful visual features, and indeed the largest asymmetry is observed in the last layer.
@@
To directly test whether joint text–image training facilitates cross-modal alignment, we performed a set of experiments with the ViT-L/14 visual encoder of the multimodal CLIP system (radford2021learning, ) on the same flickr30k dataset (Supp. Inf., Sec. J). We computed the information imbalance (II) between CLIP image representations and DeepSeek-V3 text representations and compared it with the DinoV2–DeepSeek-V3 baseline. We found that the DeepSeekV3-CLIP Image II reaches a minimum of approximately 0.3 ($\approx 15\%$ of shared features), which is lower than the minimum II observed for DinoV3-DeepSeeV3, indicating that CLIP’s visual encoder aligns more closely to the text semantics than a purely visual model of comparable capacity.
@@ 4 Limitations
It would be interesting to extend our analyses to a further variety of models of intermediate sizes and diverse architectures, including diffusion models, for example quantifying the depth and the length of the semantic regions found as a function of the number of parameters in the model, and to consider models that specifically target multilinguality euroLLM ; AyaLLM . In our experiments with translations, we used human translations from opus_books opus_books , in which the different language pairs may correspond to translations from different novels, introducing an extra source of variation not taken into account in our analysis. We explicitly avoided using LLMs to generate the translations to avoid the introduction of biases in the analyses. Indeed, we found that, for the case of English-Italian, DeepSeek-generated translations have more similar representations than those obtained from opus_books. More generally, in our multilingual analyses of DeepSeek-V3, there are more comparisons to be made, such as fixing English as a pivot language and then systematically considering different translations of the same English text into languages with different online presence, and of different degree of relatedness.

### v4

@@ 3.1.1 Heterogeneity of DeepSeek’s representations for different language pairs
We repeated the previous translation experiment with other language pairs, in particular English-Italian, English-Spanish, Italian-Spanish and German-Hungarian. We considered again sentences taken from opus books [opus_books] and applied the same filtering procedure regarding their number of tokens, as in Fig. 2a).
Fig. 2b) shows the same qualitative behavior observed for English-Spanish pairs (low Information Imbalance in the deep layers) but with major quantitative differences. Even if the training dataset of DeepSeek-V3 contains several languages in which the network can perform well, the quality of the representations is heterogeneous across languages, since it depends on the amount of data seen in each language. Fig. 2b) shows that the Information Imbalance between English and Italian is higher than between English and Spanish. A possible interpretation is that models may encode richer Spanish representations because Spanish appears much more frequently than Italian in large web corpora—Common Crawl, for instance, contains roughly twice as many Spanish as Italian documents [LanguagesUsedInternet2025, CommonCrawlLanguages2025]. Aligned with this argument, we observe that, even though Spanish and Italian are linguistically closer than either is to English, the Information Imbalance between Spanish and Italian is higher than between English-Spanish or English-Italian translation pairs. Finally, we observe that translations involving a less common language, namely Hungarian, are associated with a significantly higher Information Imbalance (the minimum value is 0.5).
Figure 3: Panel a) Minimum Information Imbalance across depth, from English to Spanish representations, as a function of the number of tokens used in the computation, for DeepSeek-V3 and Llama3.1-8b, when considering shorter (40 to 80 tokens, in circles) or longer (100 to 200 tokens, in triangles) sentences. Panels b) and c) Information Imbalance from the last token to a previous token at token-distance $\tau$, using English sentences, computed on the representations generated by DeepSeek-V3 in panel a) and by Llama3.1-8b in panel b), in different layers. Long-distance token-token correlations are maximal (Information Imbalance increases most slowly) for representations on the inner semantic layers (43 for DeepSeek-V3, 23 for Llama3.1-8b), and the effect is dramatically stronger in DeepSeek-V3. The Jackknife error bars (five repetitions subsampling 2,000 of 5,000 sentences) are too small to show; the marker sizes act as an upper bound on these errors.
@@ 3.1.2 Semantic information in LLMs is spread across many tokens
Having shown that long-span Information Imbalance captures the semantic similarity of sentences in different languages, we turn our attention to the extent to which different tokens are contributing to cross-linguistic similarity. Fig. 3a) shows the minimum Information Imbalance across all layers as a function of the number of tokens involved in the calculation, again for English-Spanish translations from opus books. The span is determined starting from the end and moving back towards the sentence beginning. For both models, we see that concatenations of tokens increases mutual predictivity: this suggests that layer summaries – for example the mean or the last tokens– do not necessarily capture all the semantic information. In facts, if we take alignment as a proxy of semantic quality, this suggests that semantic information is not concentrated in the last tokens, but spread over many of them.
Focusing first on DeepSeek-V3, we note that, up to about 10 tokens, the results do not change much between shorter (40 to 80 tokens) and longer (100 to 200 tokens) sentences. In both cases adding more tokens significantly boosts the information content. However, for shorter sentences we observe that the Information Imbalance reaches its lowest value at around $20$ tokens, and it increases with the addition of further tokens. We conjecture that this is because the last added tokens are close to the start of the sentence, and thus they do not carry much information about the semantic content of the whole sentence, given the autoregressive nature of the attention mechanism in the considered language models. In contrast, for longer sentences, we observe that adding tokens leads to a rapid decrease in Information Imbalance, which converges to very low values (around $0.05$) at around 50 tokens and stays there. On the one hand, this reflects the fact that, as longer sentences contain more information, their translation will contain more shared information at the semantic level. On the other hand, it further corroborates that information is spread across a large number of tokens. We observe similar trends for Llama3.1-8b Information Imbalance, although the curves are systematically above the DeepSeek-V3 ones. Similar to what we observed in Fig 2a), this confirms that bigger and better performing models generate representations that align better and, potentially, of better quality, capturing deep semantic similarity across different languages. This result shows that not every token contributes homogeneously to the semantic similarity between representations, and thus this analysis suggests to take the concatenation of roughly the last 20 tokens as representative of the sentence content.
@@ 3.1.3 Inner representations have longer-range correlations
The previous analysis shows that considering several tokens improves alignment across different languages One wonders if the activations of the tokens in those layers are uncorrelated or not. If activations were independent, one could argue that different tokens would be associated to different “directions” in semantic space. If they are correlated, this would suggest that the model is building a joint semantic representation of sentences spreading its information across its component tokens. We address this question by measuring how informative two distinct tokens of the same sentence are about each other. Fig. 3 (panels b) and c)) shows how much the Information Imbalance between two token positions depends on their distance, for sentences in English. Concretely, we measure $\Delta(T\rightarrow T-\tau)$, the Information Imbalance between the token at position $T$, and tokens at position $T-\tau$, for $\tau=1,2,$.... We report this measure for the first and last layers, as well as an intermediate layer from the semantic region (layers 43 of DeepSeek-V3 and 23 of Llama3.1-8b, respectively). For both LLMs, the Information Imbalance rapidly grows as a function of $\tau$ for the initial layers and saturates to 1, clearly showing that the first layer is short-range correlated. For the deeper layers that were shown in Fig. 2 to associate to the phase of shared cross-linguistic information, the Information Imbalance exhibits the slowest growth, meaning that far away tokens are predictive about each other, i.e., in these layers the information is mostly shared between tokens. DeepSeek’s deep representations have notably lower Information Imbalance across tokens than Llama3.1-8b’s, hinting at a connection between (i) the quality of the semantic representations in terms of token alignments and (ii) the magnitude and the span of long-distance correlations. A clear separation in effect strength between the intermediate semantic layer and the last layer (61 for DeepSeek-V3, 32 for Llama3.1-8b) is moreover only present for DeepSeek-V3. Still, even for DeepSeek-V3 the last layer has Information Imbalance values clearly below those of the first layer, suggesting that putatively, while the last layer is optimized to write the output, it still carries semantic content. To investigate how semantic information is distributed across tokens and whether causal asymmetries differ across languages, we conduct a detailed token-level analysis reported in Appendix I.
@@ 3.2 Images sharing high level semantics
We extend the analysis to pairs of images and measure how vision transformers (ViTs) capture semantic information for two systems: image-gpt [pmlr-v119-chen20s] and DinoV2 [oquabDINOv2LearningRobust2024]. Image-gpt performs next-token prediction after it (i) down-scales the images; (ii) unrolls them, juxtaposing the rows; (iii) quantizes the colors. It mimics the strategy of LLMs: the last layers must perform next-token prediction; previous work suggests that semantics appears in the middle pmlr-v119-chen20s. DinoV2 encodes salient features. It trains a student network to mimic its teacher and reconstruct different instances of the same image. The final layer should be the most semantically rich: its output is the input to downstream tasks such as depth estimation, image segmentation, and instance retrieval.
We first compare representations for pairs of images from the Imagenet1k dataset [Deng:etal:2009]. Each couple shares the same class, which we take as a proxy of similar semantics: we process the pairs of instances with the same architecture and compute the Information Imbalance on a random subset of 2500 same-class pairs of images, which we sample without repetition. We replicate the procedure five times to quantify uncertainty. We report the results in the left panel of Fig. 4. As the assignment of same-pair images to two spaces used to compute the Information Imbalance is arbitrary, we report the average of Information Imbalance in the two directions. The minimum of the Information Imbalance is in the middle for image-gpt and at the end for DinoV2, consistently with previous work [valeriani2023geometry]. The Information Imbalance is lower for DinoV2 ($\approx 0.43$): this would correspond to roughly 10% shared features in the Gaussian case, see Fig. 1.
@@ 3.3 Images and captions: multimodal data sharing semantic content
Image-caption pairs encode the same semantic content in different modalities: we are interested in their representations as semantics must play a key role in alignment. We process text with DeepSeek-V3 and images with DinoV2 and image-gpt. The dataset is flickr30k [youngImageDescriptionsVisual2014]: it contains pairs of images and captions; the captions are the concatenations of five different human-generated descriptions. We next repeat the analysis of Fig. 3a), estimating the minimum Information Imbalance between the image transformer and DeepSeek-V3 as a function of the number of image tokens concatenated from the end of the sequence. For DeepSeekV3, we fix the token window to 9, which corresponds to the point where the steep decrease of Information Imbalance observed in Fig. 3a) essentially saturates. We find that, for both DinoV2 and image-gpt, it is necessary to concatenate the last 200 tokens to capture the relevant shared information. Results obtained with shorter windows (last 50 tokens) and with mean-token are reported for comparison in Supp. Inf., Sec. K.
@@
In Fig. 4 we report the Information Imbalance between layer 52 of DeepSeek-V3 (captions), DinoV2 and image-gpt (images). We select layer 52 of DeepSeek-V3 because it is a layer in which the representations of text in different languages is highly mutually predictive, see Fig. 2. We recover the exact same semantic regions provided by the analysis of the image-image pairs: towards the end for DinoV2 and in the middle for image-gpt. DinoV2 has a lower Information Imbalance of 0.4. Using the Gaussian process as a reference (see Fig. 1), this would correspond to approximately 10% of shared features. The minimum Information Imbalance for image-gpt is $\approx$ 0.6, corresponding to 5% of shared features. The dashed lines in Fig. 4 report the Information Imbalance from the image to the text representation. Remarkably, for the image-gpt model the Information Imbalance is considerably larger than in the text-to-image direction. In particular, the minimum value of the Information Imbalance is $\approx 0.7$. The difference between the Information Imbalance in the two direction is significantly higher than the error bars estimated by bootstrap, indicating that the effect is statistically significant. This difference is likely due to the better quality of the DeepSeek-V3 representations. The asymmetry in the Information Imbalance is reversed for the DinoV2 model. In this case the image representation predicts the text representation marginally better than the reverse. This can be due to the fact that DinoV2 is explicitly trained to extract semantically meaningful visual features, and indeed the largest asymmetry is observed in the last layer.
@@
To directly test whether joint text–image training facilitates cross-modal alignment, we performed a set of experiments with the ViT-L/14 visual encoder of the multimodal CLIP system [radford2021learning] on the same flickr30k dataset (Supp. Inf., Sec. K). We computed the Information Imbalance between CLIP image representations and DeepSeek-V3 text representations and compared it with the DinoV2–DeepSeek-V3 baseline. We found that the DeepSeekV3-CLIP Image Information Imbalance reaches a minimum of approximately 0.3 ($\approx 15\%$ of shared features), which is lower than the minimum Information Imbalance observed for DinoV3-DeepSeeV3, indicating that CLIP’s visual encoder aligns more closely to the text semantics than a purely visual model of comparable capacity.
@@ 4 Limitations
It would be interesting to extend our analyses to a further variety of models of intermediate sizes and diverse architectures, including diffusion models, for example quantifying the depth and the length of the semantic regions found as a function of the number of parameters in the model, and to consider models that specifically target multilinguality [euroLLM, AyaLLM]. In our experiments with translations, we used human translations from opus books [opus_books], in which the different language pairs may correspond to translations from different novels, introducing an extra source of variation not taken into account in our analysis. We explicitly avoided using LLMs to generate the translations to avoid the introduction of biases in the analyses. Indeed, we found that, for the case of English-Italian, DeepSeek-generated translations have more similar representations than those obtained from opus books. We worked with a limited number of language pairs, that allowed us to spot interesting heterogeneities across representations, and to propose some plausible interpretations on why they appear. We leave for future work a detailed quantitative analysis solely dedicated to the influence of language heterogeneity in the mutual information content of representations. For example, fixing English as a pivot language and then systematically considering different translations of the same English text into languages with different online presence, and of different degree of relatedness. Finally, it would also be important to extend our analysis to embedding models trained specifically for cross-lingual retrieval, in order to assess whether architectures optimized for cross-lingual alignment exhibit different patterns of semantic convergence than general-purpose LLMs.